# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v2

@@ Appendix A Stationarity and ergodicity
Several methods have been proposed for bounding the JSR, many of which are discussed in Chang and Blondel (2013). The accompanying R package sstvars (Virolainen, 2024) implements the branch-and-bound method of Gripenberg (1996), which works well for small matrices, but can be computationally extremely demanding if the matrices are large and a tight bound is required. The JSR toolbox in MATLAB (Jungers, 2023), in turn, automatically combines various methods to substantially enhance computational efficiency.
@@ B.1 Proof of Lemma 1
We start by showing that for all $\alpha_{1,t},...,\alpha_{M,t}$, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$. First, note that by the constraints $\sum_{m=1}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$, $m=1,...,M$, we have $\alpha_{m,t}>0$ for some $m$ for all $t$. Therefore, for all $\alpha_{1,t},...,\alpha_{M,t}$, there exists invertible matrices $B_{1},...,B_{M}$ such that $B_{y,t}$ is invertible. The set of points $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ (or equally $(vec(B_{1}),...,vec(B_{M}))\in\mathbb{R}^{Md^{2}}$) for which $B_{y,t}$ is singular is $S_{t}\equiv\{(vec(B_{1}),...,vec(B_{M})\in\mathbb{R}^{Md^{2}}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$. Since the determinant of a matrix is polynomial in the entries of that matrix, and the entries of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ are linear combinations of the entries of $B_{1},...,B_{M}$, it follows that the determinant of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is a polynomial in the entries of $B_{1},...,B_{M}$. Denote this polynomial as $P_{t}(B_{1},...,B_{M})\equiv\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})$. Note that whenever $\alpha_{m,t}=0$, the impact matrices $B_{m}$ of the corresponding regimes drop out from $P_{t}(B_{1},...,B_{M})$, but in this case $P_{t}(B_{1},...,B_{M})$ can still be interpreted as a polynomial in the entries of $B_{1},...,B_{M}$ with the multipliers of the dropped entries being zero.
It then follows from Implicit function theorem that in the points $B^{0}=(vec(B_{1,0}),...,vec(B_{M,0}))\in\mathbb{R}^{Md^{2}}$ where the gradient of $P_{t}(B_{1},...,B_{M})$ is nonzero, the set $S_{t}$ is an $Md^{2}-1$-dimensional manifold locally in a neighbourhood of $B^{0}$. Because this argument applies at all points $B^{0}$ where the gradient of $P_{t}(B_{1},...,B_{M})$ is nonzero, it follows that the set $S_{t}$ is an $Md^{2}-1$-dimensional manifold in regions where this condition holds. Therefore, the Lebesque measure of the set $S_{t}$ is zero in $\mathbb{R}^{Md^{2}}$ in the region where the gradient of $P_{t}(B_{1},...,B_{M})$ is nonzero. On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $P_{t}(B_{1},...,B_{M})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$. Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$.
@@ B.2 Proof of Lemma 2
For most parts, our proof is similar to the proof of Proposition 1 in Lanne et al. (2017) making use of results from independent component analysis (Comon, 1994). First, note that from Equation (3.2), we can rewrite the model as
@@
where $\phi_{t}\equiv\sum_{m=1}^{M}\alpha_{m,t}\phi_{m,0}$ $(d\times 1)$, $\boldsymbol{A}_{t}\equiv[A_{y,t,1}:...:A_{y,t,p}]$ $(d\times dp)$, and $A_{y,t,i}\equiv\sum_{m=1}^{M}\alpha_{m,t}A_{m,i}$ contain the time-varying AR coefficients, and $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$. Then, suppose Equation (B.1) holds also for some other coefficients, impact matrix, and shocks $\tilde{\phi}_{t},\tilde{\boldsymbol{A}}_{t}$, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$, i.e., $y_{t}-\tilde{\phi}_{t}-\tilde{\boldsymbol{A}}_{t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$. By subtracting this identify from (B.1), we obtain
$\tilde{\phi}_{t}-\phi_{t}+(\tilde{\boldsymbol{A}}_{t}-\boldsymbol{A}_{t})\boldsymbol{y}_{t-1}$ $=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}$ (B.2)
@@
To complete proof, it needs to be shown that the diagonal elements of $D$ are either $1$ or $-1$. By the normalization of the unit variances of $e_{t}$ and $\tilde{e}_{t}$, $\text{Cov}(\tilde{e}_{t}|\mathcal{F}_{t-1})=\text{Cov}(Ce_{t}|\mathcal{F}_{t-1})=I_{d}$, which is equivalent to $C\text{Cov}(e_{t}|\mathcal{F}_{t-1})C^{\prime}=I_{d}$ and thereby $CC^{\prime}=I_{d}$. By substituting $C=DP$ to this expression, we obtain $DPP^{\prime}D^{\prime}=I_{d}$ and by the orthogonality of permutation matrix $DD^{\prime}=I_{d}$. That is, $d_{ii}^{2}=1$ for all $i=1,...,d$, so $d_{ii}=\pm 1$. Since diagonal elements of $D$ are, hence, either $1$ or $-1$, it follows that $B_{y,t}$ is unique up to ordering of its columns and changing all signs in a column almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.∎

### v3

@@ Appendix A Stationarity and ergodicity
Several methods have been proposed for bounding the JSR, many of which are discussed in Chang and Blondel (2013). The accompanying R package sstvars (Virolainen, 2025a) implements the branch-and-bound method of Gripenberg (1996), which works well for small matrices, but can be computationally extremely demanding if the matrices are large and a tight bound is required. The JSR toolbox in MATLAB (Jungers, 2023), in turn, automatically combines various methods to substantially enhance computational efficiency.
@@ B.1 Proof of Lemma 1
Before proving Lemma 1, we establish the following lemma that made use of in the proof:
Lemma B.1.
Suppose $P(x_{1},...,x_{k})$ is a finite-degree polynomial in the scalar variable $x_{i}\in\mathbb{R}$, $i=1,...,k$. Denote the set of vectors $(x_{1},...,x_{k})$ that satisfy the equation
$P(x_{1},...,x_{k})=0$ (B.1)
as $\mathcal{X}\equiv\{(x_{1},...,x_{k})\in\mathbb{R}^{k}|P(x_{1},...,x_{k})=0\}$. Then, the set $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{k}$, unless Equation (B.1) is identically true.
Proof.
If $k=1$, the proof is trivial. Suppose $k>1$. It then follows from Implicit function theorem that in the points $x^{0}=(x_{1,0},...,x_{k,0})\in\mathbb{R}^{k}$ where the gradient of $P(x_{1},...,x_{k})$ is nonzero, the set $\mathcal{X}$ is an $k-1$-dimensional manifold locally in a neighbourhood of $x^{0}$. Because this argument applies at all points $x^{0}$ where the gradient of $P(x_{1},...,x_{k})$ is nonzero, it follows that the set $\mathcal{X}$ is an $k-1$-dimensional manifold in regions where this condition holds. Therefore, Lebesque measure of the set $\mathcal{X}$ is zero in $\mathbb{R}^{k}$ in the region where the gradient of $P(x_{1},...,x_{k})$ is nonzero. On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $P(x_{1},...,x_{k})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{k}$. Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{k}$. ∎
To prove Lemma 1, we start by showing that for all $\alpha_{1,t},...,\alpha_{M,t}$, the matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$. First, note that by the constraints $\sum_{m=1}\alpha_{m,t}=1$ and $\alpha_{m,t}\geq 0$, $m=1,...,M$, we have $\alpha_{m,t}>0$ for some $m$ for all $t$. Therefore, for all $\alpha_{1,t},...,\alpha_{M,t}$, there exists invertible matrices $B_{1},...,B_{M}$ such that $B_{y,t}$ is invertible. The set of points $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ (or equally $(vec(B_{1}),...,vec(B_{M}))\in\mathbb{R}^{Md^{2}}$) for which $B_{y,t}$ is singular is $S_{t}\equiv\{(vec(B_{1}),...,vec(B_{M}))\in\mathbb{R}^{Md^{2}}|\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})=0\}$. Since the determinant of a matrix is polynomial in the entries of that matrix, and the entries of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ are linear combinations of the entries of $B_{1},...,B_{M}$, it follows that the determinant of $\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is a polynomial in the entries of $B_{1},...,B_{M}$. Denote this polynomial as $P_{t}(B_{1},...,B_{M})\equiv\det(\sum_{m=1}^{M}\alpha_{m,t}B_{m})$. Note that whenever $\alpha_{m,t}=0$, the impact matrices $B_{m}$ of the corresponding regimes drop out from $P_{t}(B_{1},...,B_{M})$, but in this case $P_{t}(B_{1},...,B_{M})$ can still be interpreted as a polynomial in the entries of $B_{1},...,B_{M}$ with the multipliers of the dropped entries being zero. It then follows from Lemma B.1 that $S_{t}$ has a Lebesgue measure zero in $\mathbb{R}^{Md^{2}}$.
@@ B.2 Proof of Lemma 2
The first part of our proof is largely similar to the proof of Proposition 1 in Lanne et al. (2017) making use of results from independent component analysis (Comon, 1994). First, note that from Equation (3.2), we can rewrite the model as
@@
where $\phi_{t}\equiv\sum_{m=1}^{M}\alpha_{m,t}\phi_{m,0}$ $(d\times 1)$, $\boldsymbol{A}_{t}\equiv[A_{y,t,1}:...:A_{y,t,p}]$ $(d\times dp)$, and $A_{y,t,i}\equiv\sum_{m=1}^{M}\alpha_{m,t}A_{m,i}$ contain the time-varying AR coefficients, and $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$. Then, suppose Equation (B.2) holds also for some other coefficients, transition weights, impact matrix, and shocks $\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p},\tilde{\alpha}_{m,t}$, $m=1,...,M$, $\tilde{B}_{y,t}$, and $\tilde{e}_{t}$ (and define $\tilde{\phi}_{t}$ and $\tilde{\boldsymbol{A}}_{t}$ analogously), i.e., $y_{t}-\tilde{\phi}_{t}-\tilde{\boldsymbol{A}}_{t}\boldsymbol{y}_{t-1}=\tilde{B}_{y,t}\tilde{e}_{t}$. By subtracting this identify from (B.2), we obtain
$\tilde{\phi}_{t}-\phi_{t}+(\tilde{\boldsymbol{A}}_{t}-\boldsymbol{A}_{t})\boldsymbol{y}_{t-1}=B_{y,t}e_{t}-\tilde{B}_{y,t}\tilde{e}_{t}.$ (B.3)
@@
To complete the first part of the proof, it needs to be shown that the diagonal elements of $D$ are either $1$ or $-1$. By the normalization of the unit variances of $e_{t}$ and $\tilde{e}_{t}$, $\text{Cov}(\tilde{e}_{t}|\mathcal{F}_{t-1})=\text{Cov}(Ce_{t}|\mathcal{F}_{t-1})=I_{d}$, which is equivalent to $C\text{Cov}(e_{t}|\mathcal{F}_{t-1})C^{\prime}=I_{d}$ and thereby $CC^{\prime}=I_{d}$. By substituting $C=DP$ to this expression, we obtain $DPP^{\prime}D^{\prime}=I_{d}$ and by the orthogonality of permutation matrix $DD^{\prime}=I_{d}$. That is, $d_{ii}^{2}=1$ for all $i=1,...,d$, so $d_{ii}=\pm 1$. Since diagonal elements of $D$ are, hence, either $1$ or $-1$, it follows that $B_{y,t}$ is unique up to ordering of its columns and changing all signs in a column almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
Threshold weights. To show the ”moreover”-part, consider first the threshold transition weights defined in (2.3). Denote the partition of the time periods to the regimes as $\mathcal{T}_{m}=\{t:r_{m-1}<z_{t}\leq r_{m}\}$, $m=1,...,M$, where $z_{t}$ is the switching variable and $r_{0}=-\infty,r_{M}=\infty$, and $r_{1},...,r_{M-1}$ are threshold parameters, and suppose Equation (B.2) holds also for some other partition $\tilde{\mathcal{T}}_{m}=\{t:\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}\}$. That is, we assume Equation (B.4) holds for the two sets of parameters $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $r_{1},...,r_{M-1}$ and $\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{r}_{1},...,\tilde{r}_{M-1}$.
At each $t$, the model reduces to a linear VAR $y_{t}=\phi_{0,m}+\sum_{i=1}^{p}A_{m,i}y_{t-1}+B_{m}e_{t}$ corresponding to the regime $m$ for which $r_{m-1}<z_{t}\leq r_{m}$. Consequently, Equation (B.4) can be written as
$\tilde{\phi}_{\tilde{m}_{t},0}-\phi_{m_{t},0}+\sum_{i=1}^{p}(\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i})y_{t-i}=0,$ (B.6)
where $m_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $r_{m-1}<z_{t}\leq r_{m}$ and $\tilde{m}_{t}\in\{1,...,M\}$ is the regime $m$ that satisfies $\tilde{r}_{m-1}<z_{t}\leq\tilde{r}_{m}$.
Now, Equation (B.6) defines $d$ scalar equations such that the left side of each scalar equation is a first degree polynomial in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$, where, for the $j$th equation, the coefficients are given by the $j$th element of $\tilde{\phi}_{\tilde{m}_{t},0}-\phi_{m_{t},0}$ and the elements of the $j$th rows of $\tilde{A}_{\tilde{m}_{t},i}-A_{m_{t},i}$, $i=1,...,p$. Hence, by applying Lemma B.1 to each scalar equation in (B.6), it can be concluded that the set of solutions to Equation (B.6) has a Lebesgue measure zero in $\mathbb{R}^{dp}$ unless it is identically true. Since Equation (B.6) is identically true if and only if $\tilde{\phi}_{\tilde{m}_{t},0}=\phi_{m_{t},0}$ and $\tilde{A}_{\tilde{m}_{t},i}=A_{m_{t},i}$, $i=1,...,p$, and because Equation (B.6) must hold for all $t$ (thereby for all regimes), it follows that $\tilde{\phi}_{\tilde{m},0}=\phi_{m,0}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$ almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$. It then follows from Condition (2) of Lemma 2 that once $\tilde{\phi}_{\tilde{m},0}=\phi_{m,0}$ and $\tilde{A}_{\tilde{m},i}=A_{m,i}$, $i=1,...,p$, for all $m=1,...,M$, the only way for these parameters to match the data almost everywhere is that the same regime index must apply for each $t$. Consequently, $\tilde{m}_{t}=m_{t}$ for all $t$, so that the threshold parameters $r_{1},...,r_{M-1}$ and $\tilde{r}_{1},...,\tilde{r}_{M-1}$ imply identical partitions of the data to the regimes, $\mathcal{\tilde{T}}_{m}=\mathcal{T}_{m}$ for all $m=1,...,M$ almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
Because a regime switches when (and only when) the switching variable passes one of the threshold parameter values, for given (countable number of) observations of $z_{t}$, the threshold parameters can only be identified with the resolution determined by the spacing of the order statistics of $z_{t}$. For instance, if the $n$th ($n\in\{1,...,T-1\}$) order statistic of $z_{t}$ is $1.00$, its $n+1$th order statistic is $1.02$, and $M=2$, then all values of the threshold parameter $r_{1}$ between $1.00$ and $1.02$ correspond to the same model, as they all imply the same regime-switches. Hence, the identification of the threshold parameters should be understood in the sense that they are identified to be between some order statistics of the switching variable series $z_{t}$. Now, if the threshold parameter values $\tilde{r}_{1},...,\tilde{r}_{M-1}$ are between different order statistics of $z_{t}$ than $r_{1},...,r_{M-1}$, the partitions $\mathcal{\tilde{T}}_{m}$ and $\mathcal{T}_{m}$ are different for some $m\in\{1,...,M\}$, in which case, by the above discussion, Equation (B.6) can only hold in sets of Lebesgue measure zero in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$. This concludes that the parameters $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{M-1}$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$. In other words, two distinct parameter sets give at most a measure-zero set of solutions for $(y_{1},...,y_{T})$.
Logistic weights. Next, we show the ”moreover” part for the logistic transition weights defined in (2.4) with $M=2$. Suppose Equation (B.4) holds for the two sets of parameters $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $c,\gamma$ and $\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{c},\tilde{\gamma}$. The proof proceed in the following steps: (1) we show that $\tilde{\alpha}_{m,t}=\alpha_{m,t}$ for each $t$ almost everywhere in $z_{t}\in\mathbb{R}$, (2) we use this identity to show that the parameters $c,\gamma$ are uniquely identified (almost everywhere), and then we use these results to show that (3) the intercepts and autoregression matrices are uniquely identified.
By applying Lemma B.1 to each scalar equation in Equation (B.4), it follows that Equation (B.4) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ if only if Equation (B.4) is identically true, that is, $\tilde{\phi}_{t}=\phi_{t}$ and $\tilde{\boldsymbol{A}}_{t}=\boldsymbol{A}_{t}$ (irrespective of the weight function dynamics). By making use of the identity $\alpha_{1,t}=1-\alpha_{2,t}$, these equations can be written as
$\tilde{\phi}_{1,0}-\phi_{1,0}+\alpha_{2,t}(\phi_{1,0}-\phi_{2,0})-\tilde{\alpha}_{2,t}(\tilde{\phi}_{1,0}-\tilde{\phi}_{2,0})=0$ (B.7)
and
$\tilde{A}_{1,i}-A_{1,i}+\alpha_{2,t}(A_{1,i}-A_{2,i})-\tilde{\alpha}_{2,t}(\tilde{A}_{1,i}-\tilde{A}_{2,i})=0,\ \ i=1,...,p,$ (B.8)
respectively.
Equations (B.7) and (B.8) define a system of $d(1+pd)$, $d\geq 2,p\geq 1$, linear scalar equations in the (two) variables $\alpha_{2,t}$ and $\tilde{\alpha}_{2,t}$. Therefore, for given coefficients, Equations (B.7) and (B.8) have at most one solution unless at least $d(1+pd)-1$ scalar equations are identically true. All the $d(1+pd)$ scalar equations are identically true only if $\phi_{1,0}=\phi_{2,0}$, $\tilde{\phi}_{1,0}=\tilde{\phi}_{2,0}$, $A_{1,i}=A_{2,i}$, and $\tilde{A}_{1,i}=\tilde{A}_{2,i}$ for all $i=1,...,p$, which violates Condition (2) of Lemma 2. If exactly $d(1+pd)-1$ scalar equations are identically true, the coefficients for all but one of the scalar equations are zero, and $\alpha_{2,t}$ can be expressed in the form $\alpha_{2}=a_{0}+a_{1}\tilde{\alpha}_{2,t}$ for some constants $a_{0},a_{1}\in\mathbb{R}$. Since any single solution of $\alpha_{2,t},\tilde{\alpha}_{2,t}$ can be expressed in the form $\alpha_{2,t}=a_{0}+a_{1}\tilde{\alpha}_{2,t}$ for some $a_{0},a_{1}\in\mathbb{R}$, it follows that the solution to Equations (B.7) and (B.8) must satisfy $\alpha_{2,t}=a_{0}+a_{1}\tilde{\alpha}_{2,t}$ (for all $t$).
Making the dependence of the transition weights of $z_{t}$ explicit in the notation, observe that in the identity $\alpha_{2,t}(z_{t})=a_{0}+a_{1}\tilde{\alpha}_{2,t}(z_{t})$, the logistic functions $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are both continuous and strictly increasing maps of $z_{t}$ into $(0,1)$ (as $\gamma,\tilde{\gamma}>0$). Specifically, $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ increase smoothly from 00 to $1$ in an ”S”-shaped curve as $z_{t}$ goes from $\infty$- ∞ to $\infty$. Denote $h(z_{t})\equiv\alpha_{2,t}(z_{t})-a_{0}-a_{1}\tilde{\alpha}_{2,t}(z_{t})$, which is continuous in $z_{t}$ as $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are continuous.
Suppose $h(z_{t})=0$ is not identically true, and suppose $a_{1}\neq 1$. Then, $a_{1}\tilde{\alpha}_{2,t}(z_{t})$ takes values from 00 to $a_{1}$. If $a_{1}=0$, then $h(z_{t})=0$ at most in a single point by strict monotonicity of $\alpha_{2,t}(z_{t})$. If $a_{1}<0$, then $h(z_{t})$ is strictly increasing, so $h(z_{t})=0$ at most in a single point. If $a_{1}\in(0,1)$, then $a_{1}\tilde{\alpha}_{2,t}(z_{t})<\tilde{\alpha}_{2,t}(z_{t})$, whereas if $a_{1}>1$, then $a_{1}\tilde{\alpha}_{2,t}(z_{t})>\tilde{\alpha}_{2,t}(z_{t})$ for all $z_{t}\in\mathbb{R}$. In either scenario, $h(z_{t})$ can cross zero only a finite number of times, as otherwise repeated crossing would force repeated sign changes, clearly contradicting the smooth logistic shapes that are “similar” but scaled. On the other hand, $h(z_{t})=0$ cannot hold on any set of positive Lebesgue measure because $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are strictly increasing logistic functions, and two such functions that coincide on an interval must coincide on $\mathbb{R}$, thereby forcing $a_{0}=0$ and $a_{1}=1$.
On the other hand, if $a_{1}=1$ and $a_{0}\neq 0$, then $\alpha_{2,t}(z_{t})$ and $a_{0}+a_{1}\tilde{\alpha}_{2,t}(z_{t})$ differ by constant shifting, and they can intersect in at most finitely many points. Consequently, $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ only in sets of Lebesgue measure zero in $z_{t}\in\mathbb{R}$, implying that $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ (and thereby also $\alpha_{1,t}(z_{t})=\tilde{\alpha}_{1,t}(z_{t})$) almost everywhere in $z_{t}\in\mathbb{R}$. Thus, also $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ and $\alpha_{1,t}(z_{t})=\tilde{\alpha}_{1,t}(z_{t})$ for all $t$ almost everywhere in $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
Next, we make use of the identity $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ to show that $\tilde{c}=c$ and $\tilde{\gamma}=\gamma$. By substituting the logistic weight function and simplifying, the identity $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ can be expressed in the form
$(\tilde{\gamma}-\gamma)z_{t}=\tilde{\gamma}\tilde{c}-\gamma c.$ (B.9)
If $\tilde{\gamma}=\gamma$, it follows that $\tilde{c}=c$. If $\tilde{\gamma}\neq\gamma$, Equation (B.9) has only one solution, namely, $z_{t}=(\tilde{\gamma}-\gamma)^{-1}(\tilde{\gamma}\tilde{c}-\gamma c)$. Thus, $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $z_{t}\in\mathbb{R}$.
We have established that $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$, $\tilde{\gamma}=\gamma$, and $\tilde{c}=c$, almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$. Consequently, Equations (B.7) and (B.8) simplify to
$\tilde{\phi}_{1,0}-\phi_{1,0}+\alpha_{2,t}(\phi_{1,0}-\tilde{\phi}_{1,0})+\alpha_{2,t}(\tilde{\phi}_{2,0}-\phi_{2,0})=0$ (B.10)
and
$\tilde{A}_{1,i}-A_{1,i}+\alpha_{2,t}(A_{1,i}-\tilde{A}_{1,i})+\alpha_{2,t}(\tilde{A}_{2,i}-A_{2,i})=0,\ \ i=1,...,p,$ (B.11)
respectively. Treating and $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,2$, and $c,\gamma$ as fixed constants, Equations (B.10) and (B.11) define a system of $d(1+pd)$ linear equations and $2d(1+pd)$ variables ($\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,2$). Hence, there exists at most one solution to this system linear equations. Since one solution is $\tilde{\phi}_{m,0}=\phi_{m,0},\tilde{A}_{m,1}=A_{m,1},...,\tilde{A}_{m,p}=A_{m,p}$, $m=1,2$, this is the unique solution. This concludes the proof that the parameters $\phi_{m,0},A_{m,1},...A_{m,p}$, $m=1,...,M$ as well as the parameter $r_{1},...,r_{m-1}$ models with threshold and the parameters $c,\gamma$ for models with logistics transition weights are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$.
Exogenous nonrandom weights. Suppose the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are exogenous (nonrandom). Then, arguments similar to those used for logistic weights, but applied for any finite number of regimes, show that the parameters $\phi_{m,0},A_{m,1},...,A_{M,p}$, $m=1,...,M$, are uniquely identified almost everywhere in almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.∎